OpenStudy (anonymous):

The local zoo has two water tanks for the elephant exhibit that are leaking One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h. The second water tank contains 8 gal of water and is leaking at a constant rate of 5 gal/h. When will the two tanks have the same amount of water? Explain. Let x = the number of hours the tanks are filling and let y = the number of gallons in the tank.

OpenStudy (anonymous):

There are two tanks. tank 1: y = 12 - 3*x (12 gallons and for every hour x, 3 g are lost) tank 2: y = 8 - 5*x ( 8 gallons and in every hour x, 5 g are lost)

OpenStudy (anonymous):

Okay but when will they have lost the same amount of water?

OpenStudy (anonymous):

this is the question: "When will the two tanks have the same amount of water?"

OpenStudy (anonymous):

Would you mind walking me through it I am having a hard time understanding how I am suppose to solve the equation?

OpenStudy (anonymous):

they are giving us the initial water in the tanks, and how much they lose each hour. we can construct equations that will always tell us how much is still left: total - loss per hour * hour = current volume Okay?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Okay so then for any hour we can have the water that is still in those tanks. Now,they ask us: "When will the two tanks have the same amount of water?" And the answer is: since the equations give us the amount of water in a tank then they will have the same amount of water if y1=y2

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

the independent variable is x=hours "One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h" can you construct an equation for that?

OpenStudy (anonymous):

y=12-3*x?

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

so with the other one we have those two: y=12-3*x y = 8 - 5*x

OpenStudy (anonymous):

y,,,gives us the amount of water "When will the two tanks have the same amount of water?"